Classifying a cost as either direct or indirect depends upon OA. whether the cost is expensed in the period in which it is incurred O B. whether a cost is fixed or variable ° C. whether the cost can be traced to a particular cost object in an economically feasible way O D. the behavior of the cost in response to volume changes
Answer #1

Answer:

1. C.

2. B.

3. D

4.B.

5. B.

Working (5) variable manufacturing cost = direct Material + direct manufacturing labour cost + variable manufacturing overhead

= 200 +70+65

=335

6. D.

7. B.

8. B.

9. D.

Working (9) : manufacturing cost per unit

=$2,437,500/32,500

=$75

Gross margin = (110-75) x 27,500 =$962,500

10. C.